OpenStudy (anonymous):

Consider the parabola y = 6x − x2. (a) Find the slope of the tangent line to the parabola at the point (1, 5). (b) Find an equation of the tangent line in part (a).

OpenStudy (anonymous):

((6(1+h) - (1+h)^2 ) -(6(1)-1^2) / h 6+6h -(1+2h +h^2) - 6 +1 / h 6 + 6h - 1 -2h - h^2 - 6 + 1 /h 4h - h^2 /h 4-h h ->0 so slope is 4

OpenStudy (anonymous):

y = mx + b 5 = 4(1) + b 1 = b y = 4x + 1